OpenStudy (anonymous):

Write the decimal as a fraction or a mixed number in simplest form. 3.4

OpenStudy (anonymous):

wat did u get?

OpenStudy (anonymous):

we meet again :P

OpenStudy (anonymous):

@seiga

OpenStudy (anonymous):

I got 3 2/5

OpenStudy (anonymous):

34/10=17/5

OpenStudy (anonymous):

17/5=3 2/5

OpenStudy (anonymous):

or you could say 3.4 is 3+ 0.4 so the answer will be 3 plus a fraction the fraction is 4/10 which simplifies (divide top and bottom by 2) to 2/5 so, as a mixed number, the answer is 3 2/5

OpenStudy (anonymous):

and lol @HawkCrimson

OpenStudy (anonymous):

both of u r right

OpenStudy (anonymous):

lol bye anyways(:

OpenStudy (anonymous):

yeah ur right, thats what i got lol ;p